OpenStudy (anonymous):

Write the equation of a line parallel to the line 2x + 3y = 5 and passes through the point (9, -3) A. y equals negative two-thirds x plus 3 B. y equals negative two-thirds x plus five-thirds C. y = 2x – 21 D. y equals negative two-thirds x minus 9

OpenStudy (anonymous):

medal to whoever gives best answer

OpenStudy (anonymous):

1) First find the slope from the given line. i.e., y=mx+c y=-(2/3)x+5 then m=-(2/3) 2) with that slope(m) and given point (9,-3) form the equation. i.e., y-y1=m(x-x1) y-9=-(2/3)(x+3) then y=-(2/3)x+3 option A is correct

OpenStudy (anonymous):

Tanqu...

OpenStudy (anonymous):

no thank you